OpenStudy (anonymous):

A circle has area 441 pi cm. What is the circumference?

OpenStudy (anonymous):

\[441 \pi = \pi r ^2 \rightarrow r = \sqrt{441} = 21\]

OpenStudy (anonymous):

\[r = 21cm\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

subt in 2pr (circumference/perimeter of a circle), you'll get:\[C = 2 \pi r = 42 \pi cm\]

OpenStudy (anonymous):

that's your circumference! yw :)